At the beginning of the period, assets were $200,000 and liabilities were $120,000. At the end of the period, assets were $300,000 and liabilities were $160,000. During the period, the owner withdrew $10,000 for personal use and made additional investments of $50,000. The amount of net income or net loss for the period was...?

A. Net income of $20,000

B. Net income of $40,000

C. Net income of $60,000

D. Net income of $20,000

E. Net loss of $40,000
Answer #1

Total Assets Beginning of the year = 200,000

Total Assets Ending of the year = 300,0000

Increase in Total Assets =300000-200000 = 100,000

Increase in Liabilites = 120,000-160,000 =40,000

Increase in owners Equity = 100,000-40,000 =60,000

Must be increase inonwers Equity = 60,000

Net income =

60,000= Net income -With draw +investmnet

60000 =Net inocme -10,000+50,000

= 60,000-40,000 = $20,000

Answer = Net income of $20,000